consider the ideal I=12Z of Z. Find a proper ideal J of Z such that I+J=Z
1
Expert's answer
2020-10-11T17:54:55-0400

Let us consider the ideal "J=11\mathbb Z" of "\mathbb Z". Then "1=12+(-11)\in 12\mathbb Z +11\mathbb Z=I+J". Since "I+J" is an ideal in "\mathbb Z", "a=1\cdot a\in I+J" for any "a\in\mathbb Z". Therefore, "\mathbb Z\subset I+J." Taking into account that "I+J\subset \mathbb Z", we conclude that "I+J=\mathbb Z."